# Testing the Froggatt-Nielsen Mechanism with Lepton Violation

### v1

@@ IV Results
The top-ranked textures for each mass generation mechanism are presented in Table 1 – additional textures are provided in the supplemental material of Appendix D. For both the Dirac and Majorana cases, the hierarchical structure of the charged lepton masses primarily stems from the charges of the RH charged leptons, and this fact has significant phenomenological implications. In particular, the LH rotation matrices often feature large off-diagonal terms, leading to relatively uniform contributions to various LFV processes across different textures, ultimately making the FN mechanism more predictive.
CFLV – Dirac: In most Dirac FN models, the FN scale is constrained to satisfy $\Lambda\gtrsim 10^{6}$ GeV by limits on $\mu\to e\gamma$ [42] or muon conversion in Gold [43] (see Appendix B for details). Setting $\Lambda$ to saturate its constraint within each individual model determines the highest possible rates for future CLFV signals. As shown in Fig. 1, Dirac FN textures can then predict measurable signals in LFV muon decays and $\mu$-$e$ conversion on atomic nuclei. The latter will be probed in new experiments with aluminum targets [44] (though gold targets would provide the greatest discriminatory power between textures). Furthermore, stringent constraints in the muon sector lead to suppressed LFV $\tau$ decays, with only a few textures approaching detectability in $\tau$-related channels. Overall, the predicted signal rates for FN Dirac scenarios typically exceed those of anarchic models by more than an order of magnitude, particularly for muonic processes, and offer a modest degree of discriminatory power between different textures.
@@
Type-I Seesaw: These are the most challenging scenarios to probe experimentally. All observables are governed by the scale $\Lambda$, which is fixed by the scale of neutrino masses and is too high to predict observable signals at future experiments for our best textures, with a small number of exceptions (see Figure 8 in the Appendix). However, FN type-I seesaw scenarios can be probed more effectively via $0\nu\beta\beta$ decay experiments.
@@
Figure 2: Average predicted CLFV decay rates for the top most 100 realistic Majorana FN textures (gray lines), as in Fig. 1. The top panel assumes that the scale of the Weinberg operator ($\Lambda_{W}$) coincides with the FN scale ($\Lambda$), fixing the predictions for CLFV signals from the imposed neutrino mass constraints. The bottom panel assumes $\Lambda_{W}\neq\Lambda$, such as in FN type-II seesaw scenarios, where $\Lambda$ is instead chosen to saturate its most restrictive current experimental bound of $\Lambda\sim 10^{6}$ GeV. The predicted neutrino scale for the null texture is set to $10^{14}$ GeV.
Additional information can be gained from correlations among lepton-violating observables, beyond their average predictions. For example, Fig. 3 shows how $\textrm{BR}(\mu\rightarrow 3e)$ and $\textrm{CR}(\mu\rightarrow e,\textrm{Al})$ are correlated differently for two representative FN Dirac textures. These textures were chosen because they predict very similar average values for both observables, but their widely diverging distributions makes it possible for two measurements to discriminate between these possibilities. Further correlation plots are included in the auxiliary material, see Appendix D.
Neutrino Mass Ordering – Most realistic natural textures in all scenarios favor normal ordering (NO). Inverted ordering (IO) is preferred by only a small fraction of Majorana textures and a somewhat larger subset of seesaw textures. Within the FN framework, an experimental determination of IO (e.g. by DESI measuring $\sum m_{\nu}$ with a precision of 0.02 eV [45]) would therefore strongly disfavor Dirac neutrinos, while still allowing for Majorana or type-I seesaw scenarios.
$\bm{0\nu\beta\beta}$ Decay – Predictions for $0\nu\beta\beta$ for 100 of the most realistic natural textures in Majorana and type-I seesaw models are shown in Figure 4. For both scenarios, textures with inverted ordering (IO) consistently feature values of $m_{ee}=|\sum_{i}m_{i}V^{2}_{ei}|$ at the very lower edge of the currently allowed range. We emphasize that this is not imposed on our scan and constitutes a genuine prediction of the FN mechanism – while all shown analyses impose the cosmological bound on $\sum m_{\nu}$ [46], there are no meaningful changes if we instead impose the larger laboratory bound [47]. This IO prediction lies well within the capabilities of next-generation $0\nu\beta\beta$ experiments. For textures yielding normal ordering (NO) there are significant phenomenological differences between FN Majorana and type-I seesaw. The latter remains unlikely to be detected in upcoming laboratory searches – barring a few exceptional textures – while normal-ordered FN Majorana scenarios predict $m_{ee}$ to lie well within or at most one order of magnitude below the sensitivity of upcoming laboratory searches. This opens up the tantalizing prospect of either detecting $0\nu\beta\beta$ or strongly disfavoring the entire Majorana FN framework.
@@
Figure 4: Effective Majorana mass $m_{ee}$ versus the mass of the lightest neutrino for the top 100 FN textures in the Majorana (top) and type-I seesaw (bottom) scenarios. The yellow- and blue-shaded regions indicate the allowed ranges for IO and NO, respectively, based on the current measured values for $\Delta m_{32}^{2}$ and $\Delta m_{21}^{2}$. The solid blue line indicates the strongest exclusion from current $0\nu\beta\beta$ searches, from KamLAND-Zen [46], while the dashed blue line marks the reach of the future $0\nu\beta\beta$ experiments nEXO [48] and CUPID [49]. The vertical red lines indicate the limits on $\min m_{\nu}$ in the IO and NO scenarios from current cosmological limits on $\sum m_{\nu}$. Error bars show the $1\sigma$ spread of predictions with $\delta_{\mathrm{max}}<2$.
Future Muon Colliders – Finally, we also studied collider prospects for the detection of FN-sourced lepton flavor violating processes. However, across all the realistic natural textures we found, only a small number of Dirac FN textures produces more than one $\mu\mu\to\mu\tau$ event at a future 10 TeV muon collider with 10 ab-1 of luminosity [50] (see Appendix B). It may be interesting to note that an admittedly highly speculative increase in luminosity or collider energy by one order of magnitude makes many scenarios detectable. Extending our FN setup to include the quark sector may also yield observable predictions for muon collider signals featuring both lepton- and quark-flavor violation.
@@ V Conclusions and Outlook
In this work, we have systematically explored FN models in the leptonic sector, identifying realistic natural textures for Dirac, Majorana and type-I seesaw neutrino mass generation mechanisms. Our study also sheds light on related scenarios, including type-II seesaw cases (see Appendix C.3). By extending the FN framework to leptons, we demonstrated that Dirac and Majorana FN models predict distinct correlations in detectable CLFV rates, providing characteristic signatures that set them apart from anarchic scenarios. Experimental signals are most likely to appear in the muon sector, with $\mu\rightarrow e\gamma$, $\mu\rightarrow 3e$, and $\mu$-$e$ conversion on nuclei as the most promising channels for future probes. Our results also motivate new conversion experiments in a variety of targets, most notably gold in addition to the already planned aluminum [44]. FN models with Majorana and type-I seesaw neutrinos offer testable predictions for $0\nu\beta\beta$ decay experiments, and a determination of inverted neutrino mass ordering would disfavor Dirac FN models. Proposed high-energy muon colliders fall just shy of probing purely leptonic FN scenarios. Our conclusions are unaffected by adopting either cosmological or laboratory bounds on $\sum m_{\nu}$.
@@
Note Added: While this paper was being finalized, Ref. [51] appeared on the arXiv, which explores similar topics.
Acknowledgements
@@ A.1 Identifying Realistic Natural FN Textures
This section provides details on the method used to identify natural realistic FN textures for leptons. Each texture corresponds to a specific charge assignment, and its naturalness is tested by sampling random $\mathcal{O}(1)$ coefficents for the effective Yukawa couplings. The performance of each texture is quantified by the fraction of trials yielding masses and mixings consistent with data.
@@
Permutations of fields within a family do not represent physically distinct models. To remove these redundancies, we impose the following convention for the FN charges $X_{\alpha}$, $\alpha\in\{L,e,N\}$: $|X_{\alpha_{i}}|\geq|X_{\alpha_{j}}|$ for $i<j$ when all $X_{\alpha}$ have the same sign, otherwise $X_{\alpha_{i}}\geq X_{\alpha_{j}}$. We also remove mirror charges, which are related by multiplying all charges by $-1$ and reordering them.
Each texture generates masses and mixings, which are compared to experimental values. Specifically, the observables we consider are the charged lepton masses, the PMNS matrix elements $|V_{11}|$, $|V_{13}|$, $|V_{23}|$, $|V_{21}|$, $|V_{31}|$, and $|V_{32}|$ [^1] Six PMNS elements are included because the uncertainties on these parameters are large enough that fitting only three independent elements, as would typically suffice for a unitary matrix, can still produce mixing matrices that deviate significantly from the measured values of the remaining elements., the neutrino mass squared differences, $\Delta m_{21}^{2}=m_{2}^{2}-m_{1}^{2}$ and $\Delta m_{32}^{2}=m_{3}^{2}-m_{2}^{2}$, and the cosmological bound on the sum of neutrino masses. To account for scenarios where this bound may be relaxed [52, 53, 54], we also repeated our analyses using the laboratory bound [47]. The results show no significant differences. The measured values and uncertainties for these parameters are listed in Table 2. For each observable $\mathcal{O}$, the fractional deviation from the experimental value is defined as
@@
Here $\mathcal{O}^{\rm exp}_{\rm min}$ and $\mathcal{O}^{\rm exp}_{\rm max}$ are the experimental lower and upper bounds for the observable $\mathcal{O}$, respectively. We then scan over many possible coefficient choices $c_{ij}^{\alpha}$ in Eq. (12). as explained in Section III. Textures can then be ranked by $\mathcal{F}_{x}$ for different $x$.
@@
$m_{e}$(1 TeV) $0.489535765^{+0.000000013}_{-0.000000012}$ MeV [55]
$m_{\mu}$(1 TeV) $103.3441945\pm 0.0000059$ MeV [55]
$m_{\tau}$(1 TeV) $1756.81\pm 0.16$ MeV [55]
$\Delta m_{21}^{2}$ $(7.53\pm 0.18)\times 10^{-5}\ \text{eV}^{2}$ [46]
$\Delta m_{32}^{2}$ (IO) $(-2.536\pm 0.034)\times 10^{-3}\ \text{eV}^{2}$ [46]
$\Delta m_{32}^{2}$ (NO) $(2.453\pm 0.033)\times 10^{-3}\ \text{eV}^{2}$ [46]
$\sum m_{\nu}$ (cosmo) $\leq 0.12$ eV (95% CL) [46]
$\sum m_{\nu}$ (KATRIN) $\leq 1.35$ eV (90% CL) [47]
$|V_{12}|$ $[0.513,0.579]$ [56]
$|V_{13}|$ $[0.143,0.155]$ [56]
$|V_{23}|$ $[0.637,0.776]$ [56]
$|V_{21}|$ $[0.234,0.500]$ [56]
$|V_{31}|$ $[0.271,0.525]$ [56]
$|V_{32}|$ $[0.477,0.694]$ [56]
@@ A.2 Increasingly Focused Scan Sequence
Final scans – After the secondary preselection, we perform two more scans for all mass mechanisms. First, for each remaining texture, we do a fresh scan with 1000 coefficient choices and rank textures based on $\mathcal{F}_{2}$. Textures with the same $\mathcal{F}_{2}$ are ranked by $\mathcal{F}_{5}$. Finally, for the top 1000 textures from this prescan we do a final scan with $10^{5}$ random coefficient choices. Textures are ranked by $\mathcal{F}_{x}$ with $x\approx 1.2-1.3$ depending on the FN scenario. The exact value of $x$ is chosen to be as close to 1 as possible while optimizing the statistical significance of the ranking. The resulting top 10 textures are given in Table  1.
@@
BR$(\mu^{+}\rightarrow e^{+}\gamma)$ $4.2\times 10^{-13}$ [42] $6\times 10^{-14}$ [57]
BR$(\mu^{+}\rightarrow e^{+}e^{-}e^{+})$ $1.0\times 10^{-12}$ [58] $10^{-16}$ [59]
BR$(\tau\rightarrow e\gamma)$ $3.3\times 10^{-8}$ [60] $2\times 10^{-9}$ [61]
BR$(\tau\rightarrow\mu\gamma)$ $4.2\times 10^{-8}$ [62] $10^{-9}$ [61]
BR$(\tau\rightarrow eee)$ $2.7\times 10^{-8}$ [63] $4\times 10^{-10}$ [61]
BR$(\tau\rightarrow\mu\mu\mu)$ $2.1\times 10^{-8}$ [63] $4\times 10^{-10}$ [61]
BR$(\tau^{-}\rightarrow\mu^{+}e^{-}\mu^{-})$ $2.7\times 10^{-8}$ [63] $4\times 10^{-10}$ [61]
BR$(\tau^{-}\rightarrow e^{+}\mu^{-}\mu^{-})$ $1.7\times 10^{-8}$ [63] $3\times 10^{-10}$ [61]
BR$(\tau^{-}\rightarrow e^{+}\mu^{-}e^{-})$ $1.8\times 10^{-8}$ [63] $3\times 10^{-10}$ [61]
BR$(\tau^{-}\rightarrow\mu^{+}e^{-}e^{-})$ $1.5\times 10^{-8}$ [63] $3\times 10^{-10}$ [61]
CR$(\mu^{-}\text{Ti}\rightarrow e^{-}\text{Ti})$ $6.1\times 10^{-13}$ [64] - -
CR$(\mu^{-}\text{Pb}\rightarrow e^{-}\text{Pb})$ $4.6\times 10^{-11}$ [65] - -
CR$(\mu^{-}\text{Au}\rightarrow e^{-}\text{Au})$ $7.0\times 10^{-13}$ [43] - -
CR$(\mu^{-}\text{Al}\rightarrow e^{-}\text{Al})$ $3\times 10^{-17}$ [44]
$m_{ee}$ 36 meV [46] 3 meV [48, 49]
$\mu\mu\rightarrow e\mu$
$\mu\mu\rightarrow\mu\tau$
$\mu\mu\rightarrow e\tau$
@@
Table 4: Warsaw-basis SMEFT operators included in our analysis [66]. $SU(2)_{L}$ indices are indicated by $j,k=1,2$ where necessary. Flavor indices are given by $p,r,s,t$=$1,2,3$. For leptons we include all flavor combinations. For quarks we only include flavor-diagonal combinations of the three lightest quarks, as the corresponding Wilson coefficients are enough to capture the leading contribution to $\mu-e$ conversion in nuclei. All other operators are set to zero.

### v2

@@ IV Results
The top-ranked textures for each mass generation mechanism are presented in Table 1 – additional textures are provided in the ancillary files. For both the Dirac and Majorana cases, the hierarchical structure of the charged lepton masses primarily stems from the charges of the RH charged leptons, and this fact has significant phenomenological implications. In particular, the LH rotation matrices often feature large off-diagonal terms, leading to relatively uniform contributions to various LFV processes across different textures, ultimately making the FN mechanism more predictive. Purely leptonic FN models of any type only rarely generate observable flavor-violating signals at proposed muon colliders; see Appendix B for details
CLFV – Dirac: In most Dirac FN models, the FN scale is constrained to satisfy $\Lambda\gtrsim 10^{6}$ GeV by limits on $\mu\to e\gamma$ [MEG:2016leq] or muon conversion in Gold [SINDRUMII:2006dvw] (see Appendix A.3 for details). Setting $\Lambda$ to saturate its constraint within each individual model determines the highest possible rates for future CLFV signals. As shown in Fig. 1, Dirac FN textures can then predict measurable signals in LFV muon decays and $\mu$-$e$ conversion in atomic nuclei. The latter will be probed in new experiments with aluminum targets [Mu2e:2014fns] (though gold targets would provide the greatest discriminatory power between textures). Furthermore, stringent constraints in the muon sector lead to suppressed LFV $\tau$ decays, with only a few textures approaching detectability in $\tau$-related channels. Overall, the predicted signal rates for FN Dirac scenarios typically exceed those of anarchic models by more than an order of magnitude, particularly for muonic processes, and offer a modest degree of discriminatory power between different textures.
@@
Type-I Seesaw: These are the most challenging scenarios to probe experimentally. All observables are governed by the scale $\Lambda$, which is fixed by the scale of neutrino masses and is too high to predict observable signals at future experiments for our best textures, with a small number of exceptions (see Appendix B). However, FN type-I seesaw scenarios can be probed more effectively via $0\nu\beta\beta$ decay experiments.
@@
Figure 2: Average predicted CLFV decay rates for the top most 100 realistic Majorana FN textures (gray lines), as in Fig. 1. The top panel assumes that the scale of the Weinberg operator ($\Lambda_{W}$) coincides with the FN scale ($\Lambda$), fixing the predictions for CLFV signals from the imposed neutrino mass constraints. The ‘predicted’ neutrino scale for the null texture is set to $10^{14}$ GeV. The bottom panel assumes $\Lambda_{W}\neq\Lambda$, such as in FN type-II seesaw scenarios, where $\Lambda$ is instead chosen to saturate its most restrictive current experimental bound of $\Lambda\sim 10^{6}$ GeV.
Figure 3: Effective Majorana mass $m_{ee}$ versus the mass of the lightest neutrino for the top 100 FN textures in the Majorana (top) and type-I seesaw (bottom) scenarios. The yellow- and blue-shaded regions indicate the allowed ranges for IO and NO, respectively, based on the current measured values for $\Delta m_{32}^{2}$ and $\Delta m_{21}^{2}$. The solid blue line indicates the strongest exclusion from current $0\nu\beta\beta$ searches, from KamLAND-Zen [ParticleDataGroup:2022pth], while the dashed blue line marks the reach of the future $0\nu\beta\beta$ experiments nEXO [nEXO:2021ujk] and CUPID [CUPID:2022wpt]. The vertical red lines indicate the limits on $\min m_{\nu}$ in the IO and NO scenarios from current cosmological limits on $\sum m_{\nu}$. Error bars show the $1\sigma$ spread of predictions with $\delta_{\mathrm{max}}<2$.
Additional information can be gained from correlations among lepton-violating observables, beyond their average predictions. For example, Fig. 4 shows how $\textrm{BR}(\mu\rightarrow 3e)$ and $\textrm{CR}(\mu\rightarrow e,\textrm{Al})$ are correlated differently for two representative FN Dirac textures. These textures were chosen because they predict very similar average values for both observables, but their widely diverging distributions makes it possible for two measurements to discriminate between these possibilities. Further correlation plots are included in the ancillary materials.
@@
Neutrino Mass Ordering – Most realistic natural textures in all scenarios favor normal ordering (NO). Inverted ordering (IO) is preferred by only a small fraction of Majorana textures and a somewhat larger subset of seesaw textures. Within the FN framework, an experimental determination of IO (e.g. by DESI measuring $\sum m_{\nu}$ with a precision of 0.02 eV [DESI:2016fyo]) would therefore strongly disfavor Dirac neutrinos, while still allowing for Majorana or type-I seesaw scenarios.
$\bm{0\nu\beta\beta}$ Decay – Predictions for $0\nu\beta\beta$ for 100 of the most realistic natural textures in Majorana and type-I seesaw models are shown in Figure 3. For both scenarios, textures with inverted ordering (IO) consistently feature values of $m_{ee}=|\sum_{i}m_{i}V^{2}_{ei}|$ at the very lower edge of the currently allowed range. We emphasize that this is not imposed on our scan and constitutes a genuine prediction of the FN mechanism – while all shown analyses impose the cosmological bound on $\sum m_{\nu}$ [ParticleDataGroup:2022pth], there are no meaningful changes if we instead impose the larger laboratory bound [Katrin:2024tvg]. This IO prediction lies well within the capabilities of next-generation $0\nu\beta\beta$ experiments. For textures yielding normal ordering (NO) there are significant phenomenological differences between FN Majorana and type-I seesaw. The latter remains unlikely to be detected in upcoming laboratory searches – barring a few exceptional textures – while normal-ordered FN Majorana scenarios predict $m_{ee}$ to lie well within or at most one order of magnitude below the sensitivity of upcoming laboratory searches. This opens up the tantalizing prospect of either detecting $0\nu\beta\beta$ or strongly disfavoring the entire Majorana FN framework.
@@ V Conclusions and Outlook
In this work, we have systematically explored FN models in the leptonic sector, identifying realistic natural textures for Dirac, Majorana and type-I seesaw neutrino mass generation mechanisms. Our study also sheds light on related scenarios, including type-II seesaw cases (see Appendix C.3). By extending the FN framework to leptons, we demonstrated that Dirac and Majorana FN models predict distinct correlations in detectable CLFV rates, providing characteristic signatures that set them apart from anarchic scenarios. Experimental signals are most likely to appear in the muon sector, with $\mu\rightarrow e\gamma$, $\mu\rightarrow 3e$, and $\mu$-$e$ conversion on nuclei as the most promising channels for future probes. Our results also motivate new conversion experiments in a variety of targets, most notably gold in addition to the already planned aluminum [Mu2e:2014fns]. FN models with Majorana and type-I seesaw neutrinos offer testable predictions for $0\nu\beta\beta$ decay experiments, and a determination of inverted neutrino mass ordering would disfavor Dirac FN models. Our conclusions are unaffected by adopting either cosmological or laboratory bounds on $\sum m_{\nu}$.
@@
Note Added: While this paper was being finalized, ref. [Ibe:2024cvi] appeared which takes a more literally Bayesian method to identify realistic and natural FN charge assignments for quark and lepton sectors. This analysis adopts a broadly similar approach to our work, including considering both Majorana and type-I seesaw models. Their work includes larger maximum flavor charges and considers FN in both the quark and lepton sector, which allows them to consider nucleon decay observables. On the other hand, our lepton sector FN analysis is more general in the range of allowed $\epsilon$ values, nonzero flavor charges for RH neutrinos, the inclusion of the Dirac neutrino case, and our predictions for CLFV observables. Our analyses are therefore highly complementary.
Acknowledgments
@@ A.1 Identifying Realistic Natural FN Textures
This section provides details on the method used to identify natural realistic FN textures for leptons.
@@
Permutations of fields within a family do not represent physically distinct models. To remove these redundancies, we impose the following convention for the FN charges $X_{\alpha}$, $\alpha\in\{L,e,N\}$: $|X_{\alpha_{i}}|\geq|X_{\alpha_{j}}|$ for $i<j$ when all $X_{\alpha}$ have the same sign, otherwise $X_{\alpha_{i}}\geq X_{\alpha_{j}}$. We also remove mirror charges, which are related by multiplying all charges by 1-1 and reordering them.
Each texture generates masses and mixings, which are compared to experimental values. Specifically, the observables we consider are the charged lepton masses, the PMNS matrix elements $|V_{11}|$, $|V_{13}|$, $|V_{23}|$, $|V_{21}|$, $|V_{31}|$, and $|V_{32}|$ [^2] Six PMNS elements are included because the uncertainties on these parameters are large enough that fitting only three independent elements, as would typically suffice for a unitary matrix, can still produce mixing matrices that deviate significantly from the measured values of the remaining elements., the neutrino mass squared differences, $\Delta m_{21}^{2}=m_{2}^{2}-m_{1}^{2}$ and $\Delta m_{32}^{2}=m_{3}^{2}-m_{2}^{2}$, and the cosmological bound on the sum of neutrino masses. To account for scenarios where this bound may be relaxed [Oldengott:2019lke, Chacko:2019nej, Escudero:2022gez], we also repeated our analyses using the laboratory bound [Katrin:2024tvg]. The results show no significant differences. The measured values and uncertainties for these parameters are listed in Table 2. For each observable $\mathcal{O}$, the fractional deviation from the experimental value is defined as
@@
Here $\mathcal{O}^{\rm exp}_{\rm min}$ and $\mathcal{O}^{\rm exp}_{\rm max}$ are the experimental lower and upper bounds for the observable $\mathcal{O}$, respectively. We then scan over many possible coefficient choices $c_{ij}^{\alpha}$ as explained in the main text. Textures can then be ranked by $\mathcal{F}_{x}$ for different $x$.
@@
$m_{e}$(1 TeV) $0.489535765^{+0.000000013}_{-0.000000012}$ MeV [Xing:2011aa]
$m_{\mu}$(1 TeV) $103.3441945\pm 0.0000059$ MeV [Xing:2011aa]
$m_{\tau}$(1 TeV) $1756.81\pm 0.16$ MeV [Xing:2011aa]
$\Delta m_{21}^{2}$ $(7.53\pm 0.18)\times 10^{-5}\ \text{eV}^{2}$ [ParticleDataGroup:2022pth]
$\Delta m_{32}^{2}$ (IO) $(-2.536\pm 0.034)\times 10^{-3}\ \text{eV}^{2}$ [ParticleDataGroup:2022pth]
$\Delta m_{32}^{2}$ (NO) $(2.453\pm 0.033)\times 10^{-3}\ \text{eV}^{2}$ [ParticleDataGroup:2022pth]
$\sum m_{\nu}$ (cosmo) $\leq 0.12$ eV (95% CL) [ParticleDataGroup:2022pth]
$\sum m_{\nu}$ (KATRIN) $\leq 1.35$ eV (90% CL) [Katrin:2024tvg]
$|V_{12}|$ $[0.513,0.579]$ [Esteban:2020cvm]
$|V_{13}|$ $[0.143,0.155]$ [Esteban:2020cvm]
$|V_{23}|$ $[0.637,0.776]$ [Esteban:2020cvm]
$|V_{21}|$ $[0.234,0.500]$ [Esteban:2020cvm]
$|V_{31}|$ $[0.271,0.525]$ [Esteban:2020cvm]
$|V_{32}|$ $[0.477,0.694]$ [Esteban:2020cvm]
@@ A.2 Increasingly Focused Scan Sequence
Final scans – After the secondary preselection, we perform two more scans for all mass mechanisms. First, for each remaining texture, we do a fresh scan with 1000 coefficient choices and rank textures based on $\mathcal{F}_{2}$. Textures with the same $\mathcal{F}_{2}$ are ranked by $\mathcal{F}_{5}$. Finally, for the top 1000 textures from this prescan we do a final scan with $10^{5}$ random coefficient choices. Textures are ranked by $\mathcal{F}_{x}$ with $x\approx 1.2-1.3$ depending on the FN scenario. The exact value of $x$ is chosen to be as close to 1 as possible while optimizing the statistical significance of the ranking.
@@
[…]
$m_{ee}$ 36 meV [ParticleDataGroup:2022pth] 3 meV [nEXO:2021ujk, CUPID:2022wpt]
$\mu\mu\rightarrow e\mu$ - - - -
$\mu\mu\rightarrow\mu\tau$ - - - -
$\mu\mu\rightarrow e\tau$ - - - -
@@
Table 4: Warsaw-basis SMEFT operators included in our analysis [Grzadkowski:2010es]. $SU(2)_{L}$ indices are indicated by $j,k=1,2$ where necessary. Flavor indices are given by p,r,s,tp,r,s,t=$1,2,3$. For leptons we include all flavor combinations. For quarks we only include flavor-diagonal combinations of the three lightest quarks, as the corresponding Wilson coefficients are enough to capture the leading contribution to $\mu-e$ conversion in nuclei. All other operators are set to zero.